OpenStudy (anonymous):

Can someone tell me if I answered this question correctly? A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 2 experience points. On level 2, a player must earn 4 experience points. On level 3, a player must earn 7 experience points. On level 4, a player must earn 11 experience points. Which recursive equation represents the pattern? A. an = an – 1 + n B. an = an – 1 – n C. an = an – 1 + 2n --> My answer D. an = an – 1 – 2n

OpenStudy (anonymous):

@phi Can you please check this for me.

OpenStudy (anonymous):

@satellite73

OpenStudy (phi):

I would say \[ a_n= a_{n-1} + n \] for example, \[ a_2 = a_1 +2+1 \\ a_2 = 2 + 2 = 4 \] and \[ a_3 = a_2 + 3 \\ a_3= 4 + 3 = 7 \]